One side of triangle is 3 times the shortest side. The third side is 19 feet more than the shortest side. The perimeter is 94 . Find all three sides.
Answer by JulietG(1812)

P = A + B + C
P = 94
C = B+19 (B is the shortest side)
A = 3B
Substitute the known values of A and C from the bottom equations and plug them into the top one.
P = A + B + C
94 = A + B + C
94 = (3B) + B + (B+19)
94 = 5B + 19
Subtract 19 from each side
75 = 5B
Divide each side by 5
15 = B
If the shortest side is 15, then side A is 45, and side C is 34
A = 45
B = 15
C = 34
45 + 15 + 34 = 94